Heparin 70,000 units in 500 mL D5W - run cont @ 5800 units/hr. How many ml/hr will you need to set your pump for?
Answer #1

ANSWER.

* Calculation of ml/hr.

           ml/hr = Volume(ml) / Time(hr).

Calculation of volume.

     Volume(ml) = Desired dose(D) / Dose in hand(H) * Quantity(Q).

Here

    Desired dose(D) = 5800 unit.

   Dose in hand(H) = 70000 unit.

      Quantity(Q) = 500 ml.

As per above formula

      Volume(ml) of drug required = D/H * Q

                                                    = 5800/70000 * 500

                                                    = 0.083 * 500

                                                    = 41.5 ml.

                        Time(hr) = 1 hr.

        So

                 ml/hr = 41.5 ml/hr.